OpenStudy (albert0898):

In the equation x^2 + kx + 54 = 0, one root is twice the other root. The value(s) of k is (are)... (A) -5.2 (B) 15.6 (C) 22.0 (D) +- 5.2 (E) +- 15.6

OpenStudy (albert0898):

\[x^2 + kx + 54 = 0\]

OpenStudy (roadjester):

So, you want one root to be twice the other. You'll need something along the lines of \[(x+a)(x+2a)\] Now foil it to get \[x^2+3ax+2a^2=x^2+kx+54\]

OpenStudy (albert0898):

What do I do after that? Does that mean that 3ax = kx and that 2a^2 = 54

OpenStudy (roadjester):

Yes, and you now have a system of equations.

OpenStudy (albert0898):

I get it now. Thank you so much! You cleared everything up for me